### v1

# Statistical Demography Meets Ministry of Health: The Case of the Family Planning Estimation Tool

Abstract
The Family Planning Estimation Tool (FPET) is used in low- and middle-income countries to produce estimates and short-term forecasts of family planning indicators, such as modern contraceptive use and unmet need for contraceptives. Estimates are obtained via a Bayesian statistical model that is fitted to country-specific data from surveys and service statistics data. The model has evolved over the last decade based on user inputs.
This paper summarizes the main features of the statistical model used in FPET and introduces recent updates related to capturing contraceptive transitions, fitting to survey data that may be error prone, and the use of service statistics data. We use our experience with FPET to discuss lessons learned and open challenges related to the broader field of statistical modeling for monitoring of demographic and global health indicators to help further optimize the application of statistical modeling in practice.
1 Introduction
The elevation of family planning (FP) on the global stage in 2012 with the launch of FP2020 provided an opportunity to create new approaches for monitoring of FP programs. The need for annual estimates to track progress of the initiative led to the development of standard indicators, approaches, and methodologies. For countries, having annual estimates of FP indicators was a priority to better gauge progress towards their own objectives. After the FP2020 initiative concluded, the FP2030 initiative was started with a continued focus on measurement and setting of country-specific commitments.
Track20, implemented by Avenir Health and funded by the Gates Foundation, is a global family planning project aimed at improving global and country level use of data. The project’s aims include the development and introduction of innovative and user-friendly methodologies, tools, and approaches that build capacity and broaden who can effectively engage with family planning data. Track20 has provided the annual estimates used for FP2020 related measurements and continues to assess progress for FP2030 and the Ouagadougou Partnership (focused on FP in countries in western Africa, see https://www.speakupafrica.org/program/speak-up-africa-website-blurb-for-the-ouagadougou-partnership-op/). This is done through a bottom-up reporting process led by government monitoring and evaluation (M&E) officers, effectively linking country and global FP data.
The Family Planning Estimation Tool (FPET) is used by Track20 in low- and middle-income countries to produce estimates and short-term forecasts of FP indicators, such as modern contraceptive use and unmet need for contraceptives. Estimates are obtained via a Bayesian statistical model that is fitted to country-specific data from surveys and service statistics data. The model is a country-specific implementation of a global model for FP estimation, referred to as the global Family Planning Estimation Model (Alkema et al. [2013], Cahill et al. [2018], Kantorová et al. [2020]). The country-specific implementation was produced to introduce and facilitate in-country usage by the Track20 project. FPET has evolved over the last decade based on usage in the Track20 project and other user inputs and requests. Recent statistical modeling updates include relaxation of strong parametric assumptions made regarding contraceptive use transitions (Susmann and Alkema [2023]), updated use of survey data to improve estimation in the presence of data outliers (Alkema et al. [2024a]), and updated use of service statistics data to better account for uncertainty associated with these data (Mooney et al. [2024a]).
In this paper, we present the current version of the Family Planning Estimation Model and Tool. We first introduce the how FPET came about, how it is used in the Track20 project, and how the modeling approach has evolved. The methods section presents the current model, highlighting how different advances in statistical modeling were incorporated. We use our experience with FPET to discuss lessons learned and open challenges related to the broader field of statistical modeling for monitoring demographic and global health indicators.
2 Background
2.1 Why a model is needed to produce FP estimates
Suppose that we are interested in the proportion of married or in-union women aged 15-49 within a population who are using (or whose partner is using) a modern contraceptive method, abbreviated as mCPR. Data on mCPR can be obtained from household surveys in which participants report whether they or their partner are currently using a modern contraceptive method. Such household surveys may be conducted by international organizations or by local organizations or governments. Survey programs and categorizations include the Demographic and Health Survey Program (DHS), Performance Monitoring for Action (PMA), UNICEF Multiple Indicator Cluster Surveys (MICS), national surveys, and other surveys. Micro data from each survey can be used to calculate estimates of mCPR for the population-time period covered by the survey, taking into account the complex sampling design of each survey, as well as a measure of the uncertainty in the mCPR estimate owed to the survey design, referred to as sampling error.
Often survey data alone for some indicator of interest may not be sufficient to inform programs and understand progress. First, data are not necessarily available for all years of interest, e.g., for years in the past, the period after the most recent data point until the current year, and future years. Moreover, data are subject to measurement error, which may be substantial. This issue is discussed in detail in Alkema et al. [2024a]. An example is shown in Figure 1 for Burundi. In Burundi, the most recent national survey, from 2012, suggests a large fluctuation in mCPR which is likely to be due to measurement error. This example illustrates the need to integrate data sources to produce reliable estimates and forecasts.
Figure 1: Data for mCPR among married women for Burundi. Data are shown in colored dots with 95% confidence intervals that reflect sampling errors.
2.2 What FPET does
FPET produces estimates of FP indicators using available data for a population of interest, see Figure 2. In FPET, the FP indicators considered are contraceptive use and unmet need for contraceptives. Methods are categorized into modern versus traditional methods. Modern methods of contraception include female and male sterilization, oral hormonal pills, the intra-uterine device, male and female condoms, injectables, the implant (including Norplant; Wyeth-Ayerst, Collegeville, PA, USA), vaginal barrier methods, standard days method, lactational amenorrhea method, and emergency contraception. Traditional methods of contraception include abstinence, the withdrawal method, the rhythm method, douching, and folk methods. Unmet need for family planning is defined as the percentage of women who want to stop or delay childbearing but who are not currently using any method of contraception to prevent pregnancy. Also included are women who are currently pregnant or postpartum amenorrheic whose pregnancies were mistimed or unwanted.
Figure 2: The Family Planning Estimation Tool (FPET) produces estimates of FP indicators, based on survey data and observations on estimated modern use (EMUs), obtained from service statistics. The graph shows survey data on modern contraceptive use for married women of reproductive age in Kenya, together with an FPET fit (point estimates are given by the solid black line, shaded areas represent 90% credible intervals).
FP indicators are estimated for women aged 15 to 49 years old, considering women’s marital status. More precisely, estimates are produced separately for women who are married or in a union (referred to as married women in this paper), versus women who are not married. Sexual activity is considered in the calculation of the indicators, as summarized in Kantorová et al. [2020].
Data on FP are available from household surveys and routine data collection. As mentioned in the previous section, contraceptive prevalence data obtained from surveys refers to the percentage of women who report themselves or their partners as currently using at least one contraceptive method of any type (modern or traditional). Service statistics data refer to data obtained from routine data collection. These systems may record various FP-related outcomes, including FP commodities distributed to clients or facilities, visits to FP facilities/providers, or family planning users. Track20 has developed a tool that converts service statistics data into a single metric representing volume of services, Estimated Modern Use (EMU), discussed further in a later section.
FPET produces estimates of FP indicators using available survey and service statistics data for a population (e.g., a country or subnational region). An illustration is given in Figure 2 for Kenya.
3 Usage of FPET in the Track20 project
3.1 Overview of Track20’s approach
The Track20 project works to build the capacity of countries to generate and use data to inform their programming, improve impact, and accelerate progress toward their FP goals. One of the pillars of the Track20 approach is the cultivation of a network of M&E Officers dedicated to increasing the quality and use of FP data. These M&E Officers work within or in partnership with Ministries of Health in 34 countries. This approach positions governments to be the driving force of FP data instead of relying on implementing partners to produce needed analyses and findings. Track20 provides ongoing on-demand technical support to these countries. This includes an annual training that brings all M&E Officers together to learn additional analytical skills, engage in south-to-south learning, and to identify opportunities for new analyses and tools based on current government challenges. This direct exchange with multiple countries enables Track20 to develop tools that directly speak to current government priorities. It also provides an opportunity for any new tools to be vetted and tried across many countries, providing valuable feedback that improves Track20’s work.
Track20 M&E officers use FPET to assess progress towards country strategies and commitments. The officers are trained in FPET and use it to produce FP estimates for their country. These FPET estimates are reviewed, vetted, and approved through annual consensus meetings. These meetings are convened by governments and attended by partners and donors, providing a time to reflect on progress and identify any barriers or opportunities moving forward.
3.2 Evolution of FPET
The estimation of FP indicators in FPET has evolved based on user inputs and further model developments. The first version was based on a global modeling approach to estimate and project FP indicators for married women aged 15-49 in all countries in the world using survey data (Alkema et al. [2013]). For in-country usage in the Track20 project, we produced FPET based on a local version of the global model, allowing users to fit the model to data from one country only (New et al. [2017]). This version expanded upon the global model by allowing for the use of service statistics data.
Subsequently, FPET evolved to better capture local contexts. Model updates were introduced to improve predictive performance and to better account for survey data quality issues (Cahill et al. [2018]), and to improve the use of service statistics data (Cahill et al. [2021]). The population considered was extended from married women to all women, using a local implementation of the global model for estimation among unmarried women (Kantorová et al. [2020], Guranich et al. [2021]). In 2024, several updates were introduced to further improve FPET’s utility. The next section introduces the FPET model specification based on these updates.
4 FPET statistical model specification and selected findings
4.1 Overview
We start by introducing notation to formalize the problem of producing estimates and forecasts of FP indicators. We are interested in the proportion of women (by marital status) who either use a modern method of contraception, use a traditional method, have an unmet need for any method, or do not have a demand for FP (no need group). We denote these proportions as $\phi_{j}$ for $j=1,\ldots,4$ referring to modern, trad, unmet any, and no need respectively. Unmet need for modern methods is defined as the sum of traditional users and women with an unmet need for any method. Demand $\eta_{1}$ is defined as the sum of modern users, traditional users, and unmet need for any method: $\eta_{1}=\phi_{1}+\phi_{2}+\phi_{3}$. Demand satisfied is defined as the modern users over demand, $\eta_{2}=\phi_{1}/(\phi_{1}+\phi_{2}+\phi_{3})$. When referring to specific population-periods, additional indices are used. In particular, $\phi_{j,c,m,t}$ refers to the $j$-th proportion for country $c$, marital group $m$, and year $t$.
To produce estimates and forecasts for FP indicators, we made assumptions in the global model about (1) how FP indicators may change over time and vary across different populations, and (2) how available data relate to the indicators. In model descriptions, we distinguish between these two sets of assumptions and refer to the first set as the process model and the second set as the data model assumptions. We introduce the process and data models for modern use and unmet need for modern methods below. The appendix includes the specification of the process and data models for traditional use.
4.2 Process model: How do FP indicators vary over time and differ across populations?
Capturing contraceptive transitions
In the process model, we consider how FP indicators change over time. The main goal in FPET is to capture long-term and short-term changes. We do so using a Bayesian hierarchical B-spline transition model with autoregressive country-deviations terms, equivalent to an ARIMA(1,1,0) model with time-varying drift.
Over longer time periods, the main assumptions underlying FPET are based on capturing a contraceptive transition, introduced by Alkema et al. [2013], Cahill et al. [2018] and Susmann and Alkema [2023]. Specifically, over longer time periods, we assume that demand for contraceptives increases from low to high levels, where increases start slowly at low demand, accelerate once growth has been initiated, and then rates of change slow down as demand reaches high levels, which we refer to as an asymptote. Similarly, we assume that demand satisfied with modern methods also follows this type of increase. Illustrations of observed trends globally, that support these assumptions, are given in Figure 3. Based on combining the two transitions, we obtain modern contraceptive use (mCPR) and unmet need for modern methods. Resulting estimates show a typical pattern in unmet need whereby unmet need increases at the start of a transition, when supply is lagging demand, followed by a decrease once supply catches up with demand.
Figure 3: Observed data on demand and demand satisfied for modern methods for selected countries over time. Observations refer to prevalence among married women of reproductive age. Countries are included in the selected main regions if they have more than 4 observations.
The increase in demand $\eta_{1,c,m,t}$, and demand satisfied with modern methods $\eta_{2,c,m,t}$, from low levels to an asymptote are modeled with flexible transition models (Susmann and Alkema [2023]). In this type of model, the rate of change in an indicator of interest $\eta_{k,c,m,t}$ is parameterized using a systematic component and a smoothing component, where the systematic component captures the expected long-term transition from low to high levels of the indicator while the local smoothing component captures deceleration or acceleration as indicated by the data. Specifically, and dropping indices $k$ and $m$ for notation convenience, a transition model for indicator of interest $\eta_{c,t}$ is specified as follows:
$\Delta\text{logit}(\eta_{c,t})=f(\eta_{c,t-1},\bm{\beta}_{c},\lambda_{c})+\varepsilon_{c,t},$ (1)
where the transition function $f(\eta_{c,t-1},\bm{\beta}_{c},\lambda_{c})$ captures the expected rate of change of the logit-transformed indicator and the stochastic smoothing component $\varepsilon_{c,t}$ captures deviations away from the expected rate of change. The specification for $\eta_{c,t}$ is completed through defining its level in a reference year $t^{*}$:
$\text{logit}(\eta_{c,t^{*}})=\Omega_{c}.$ (2)
The specification of the transition function $f(\eta_{c,t-1},\bm{\beta}_{c},\lambda_{c})$ follows that of Susmann and Alkema [2023], see Figure 4. The transition function is parameterized using B-splines, as follows:
$f(\eta_{c,t-1},\bm{\beta}_{c},\lambda_{c})=\sum_{j=1}^{J}\bm{\beta}_{c,j}B_{j}\left(\eta_{c,t}/\lambda_{c}\right),$ (3)
where $B_{j}(\cdot)$ refers to the $j$th spline basis function, $\bm{\beta}_{c}$ is the vector of splines parameters, and $\lambda_{c}$ is the asymptote. An illustration of this transition function is given in Figure 4.
Figure 4: Illustration of the transition model specification for indicator $\eta_{c,t}$. Left: Rate of change (black) as a function of the $\eta_{c,t}$, parameterized using B-splines. The gray lines illustrate a scaled version of the non-zero spline basis functions. Right: Transition of $\eta_{c,t}$ from low to high levels, based on the rate of change from plot on left. Figure adapted from Susmann and Alkema [2023].
The smoothing terms $\varepsilon_{c,t}$ represent deviations from the smooth transition functions, and are incorporated to capture short-term accelerations or stalls as suggested by the data. The smoothing terms are assigned an AR(1) model:
$\varepsilon_{c,t}|\varepsilon_{c,t-1},\rho^{(AR)},\sigma_{\varepsilon}\sim N(\rho^{(AR)}\cdot\varepsilon_{c,t-1},\sigma_{\varepsilon}^{2}),$
with unconditional standard deviation $\text{sd}(\varepsilon_{c,t})=\sigma_{\varepsilon}/\sqrt{1-\rho^{(AR)2}}$. Using an AR(1) smoothing term in the expression for the rate of change implies that the model for $\eta_{c,t}$ itself can be seen as an ARIMA(1,1,0) model on the logit scale with time-varying drift term $f(\eta_{c,t-1},\bm{\beta}_{c},\lambda_{c})$.
Sharing information across populations
Our goal is to estimate and forecast FP indicators from 1970 until 2030 for all countries globally, for women of reproductive age, separately by marital/in-union status. However, a country’s data may not be sufficient to estimate FP for the time period of interest. For example, for countries that are at the start of their transition, country data alone do not allow for estimating its asymptotes. Some countries may have very limited data on FP among unmarried women.
To overcome data limitations, we fit a global model in which information is shared across populations using hierarchical models. We fit the model separately for married versus unmarried women, given that transitions differ between these groups. For each model, we group countries based on geo-political characteristics into clusters and subclusters of countries, following the groupings proposed by Bearak et al. [2020]. In addition, for the modeling of FP among unmarried women, we further extend the groupings to account for differences in cultural norms related to FP and sexual activity among unmarried women, following the approach by Kantorová et al. [2020].
The hierarchical models used, for each marital group, are as follows. Three sets of model parameters are given hierarchical priors:
  1. 1.
    asymptotes $\lambda_{k,c}$ for $k=1,2$ controlling the limiting value of an indicator;
  2. 2.
    spline coefficients $\bm{\beta}_{k,c,j}$ for $k=1,2$ determining the transition function $f$; and
  3. 3.
    $\Omega_{k,c}$ for $k=1,2$, the level of an indicator in a reference year
Specifically, hierarchical priors are assigned to transformations of the model parameters, to constrain parameters to appropriate ranges.
Let $\theta_{\gamma}^{(l)}$ denote the value of one of the transformed parameters for a geographic unit indexed by $\gamma$ at hierarchical level $l$. Level $l$ ranges from $l=1$ referring to the global level, to $l=2$ cluster, $l=3$ subcluster, and $l=4$ country-level. We adopt the following setup:
$\theta_{\gamma}^{(l)}\mid\theta^{(l-1)}_{\gamma^{\prime}},\sigma^{(l)}_{\theta}$ $\sim N(\theta^{(l-1)}_{\gamma^{\prime}},\sigma^{(l)}_{\theta}),\text{ for }l>1,l>1.$
Here, $\theta^{(l-1)}_{\gamma^{\prime}}$ is the parameter value at one level higher in the hierarchy for the region $\gamma^{\prime}$ that contains unit $\gamma$. At the top level of the hierarchy, $l=1$, only a single parameter is estimated, representing an intercept that applies to all lower level geographic units. The parameter $\sigma_{\theta}^{(l)}$ for $l>1$ describes variability of values of the parameter $\theta$ across geographic units at level $l$ of the hierarchy.
Illustrative findings
Illustrative results of model fits are shown in Figure 5 for Angola, Bolivia, and Burundi. We see that the model can produce estimates and forecasts for all years. Uncertainty associated with FP indicators increases with the forecast horizon.
Figure 5: Data, estimates, and forecasts from 1970 until 2030 for FP indicators for married women in Angola, Bolivia, and Burundi. Graphs show mCPR, unmet need for modern methods, demand, and demand satisfied with modern methods. Survey data are shown in colored dots with 95% confidence intervals that reflect total error variance. FPET estimates are shown in black, with grey shaded areas representing 90% credible intervals.
4.3 Data model for survey data
The model is fitted to survey data on FP indicators, while accounting for different sources of error. First, all observations are subject to sampling error arising from the survey design. In addition, other non-systematic errors may occur which we will refer to as measurement errors. Measurement errors may occur due to errors introduced during data collection or compilation. The extent of such errors may vary based on survey characteristics. For example for mCPR estimation, some survey programs have been found to provide more variable measurements of mCPR than others (Alkema et al. [2013]). Extreme errors may also occur, as we saw in the mCPR example for Burundi in Figure 1, resulting in outlying observations due to measurement error. Finally, errors may occur if the sampled population differs from the population of interest. For mCPR estimation, such errors are typically absent from recently collected data by international programs but may be present in older data or data from national or other survey programs. Examples include variation in age groups (Alkema et al. [2013], Kantorová et al. [2020]). When using survey data, we need to account for data errors, to produce estimates that are aligned with observations that are very certain and deemed to be close to the true value, while smoothing over outlying data points that are very uncertain.
The survey data model specifies the relation between the observed survey data and the FP indicators that are estimated. In FPET, the normal-with-optional-shrinkage (NOS) class of data models is used to account for the types of errors that may occur (Alkema et al. [2024a]). This class of data models accounts for sampling errors and differences in observational uncertainty based on survey characteristics. In addition, the NOS data model uses shrinkage priors to produce estimates that are robust to outlying observations. The data model class also accounts for settings where errors may be correlated across observations, as is the case for FP with observations from PMA data series.
Details on the the NOS data model and its use in FPET are provided in Alkema et al. [2024a]; we summarize the main set up here. We consider the NOS data model for the logit-transform of two observed proportions: mCPR and unmet need for modern methods out of non-modern-users. With slight misuse of notation for ease of exposition, let $y_{i}$ refer to one observed proportion and let $\phi_{c[i],t[i]}$ refer to the true latent indicator for the corresponding country-year. For example, let $y_{i}$ refer to observed mCPR and $\phi_{c[i],t[i]}$ to true mCPR. The total error associated with $y_{i}$, on the logit-transformed scale, is decomposed into type-specific errors:
$E_{i}$ $=\mathrm{logit}(y_{i})-\mathrm{logit}(\phi_{c[i],t[i]}),$ (4)
$=E_{i}^{(\mathsf{samp})}+\mathbb{I}(d[i]\in\mathcal{D})E_{i}^{(\mathsf{source})}+\mathbb{I}(i\in\mathcal{C})E_{i}^{(\mathsf{char})}+\mathbb{I}(i\in\mathcal{I})E_{i}^{(\mathsf{outlier})},$ (5)
where sampling error is given by $E_{i}^{(\mathsf{samp})}$, source-type specific measurement error by $E_{i}^{(\mathsf{source})}$, $E_{i}^{(\mathsf{char})}$ captures differences if the sampled population differs from the population of interest, and finally, observation-specific error for outliers is given by $E_{i}^{(\mathsf{outlier})}$. Indicator functions indicate whether the error term is assigned to each observation.
In the NOS data model applied to FP indicators, each type-specific error term is assumed to be normally distributed with mean zero and standard deviation $\sigma_{i}^{(\mathsf{type})}$:
$E_{i}^{(\mathsf{type})}|\sigma_{i}^{(\mathsf{type})}$ $\sim N\left(0,\sigma_{i}^{(\mathsf{type})2}\right).$ (6)
For the sampling error $E_{i}^{(\mathsf{samp})}$, the associated variance is $\sigma_{i}^{(\mathsf{samp})2}=s_{i}^{2}$, where $s_{i}$ is the sampling error on the logit-scale. Source-type errors capture that errors from some programs may be larger than errors from other programs: $\sigma_{i}^{(\mathsf{source})}=\tilde{\sigma}_{d[i]}^{(\mathsf{source})}$, where $d[i]\in\{1,\dots,D\}$ indexes the survey source of observation $i$, i.e., DHS, MICS, PMA, National survey, or other survey. For mCPR estimation, we assign source errors to observations from all sources except those from DHS, i.e., $\mathcal{D}$ excludes DHS. The term $E_{i}^{(\mathsf{char})}$ with variance $\sigma^{(\mathsf{char})2}$ is added if a survey did not capture the entire population of interest, for example, if ever-married women were interviewed as compared to married women. The subset of observation indices for which the population surveyed differs from that of interest, $\mathcal{C}$, is obtained from survey meta data.
The outlier error variance arises from a shrinkage prior on the error term (Piironen and Vehtari [2017]) that shrinks outlier errors towards zero for most observations, only allowing large errors when the observation is extremely different from the trend. In this set up, the standard deviation term $\sigma_{i}^{(\mathsf{outlier})}$ is parameterized using a scale parameter $\tau$ that is shared across observations, and an observation-specific scale parameter $\gamma_{i}$ that allows individual errors to escape from shrinkage, $\sigma_{i}^{(\mathsf{outlier})}=\tau\tilde{\gamma}_{i}=\sqrt{\frac{\tau^{2}\vartheta^{2}\gamma_{i}^{2}}{\vartheta^{2}+\tau^{2}\gamma_{i}^{2}}}$. The local scale parameter $\gamma_{i}$, which is assigned a heavy-tailed half-Cauchy prior, $\gamma_{i}\sim C^{+}(0,1)$. Observation indices $\mathcal{I}$ refer to the set of data points that may be possibly outlying, determined by a data classification algorithm [Alkema et al., 2024a].
Based on the expression for the error terms in Eq. (6), and assuming independence across errors for a single observation, the marginal density for individual observations is given by
$\mathrm{logit}(y_{i})|\phi_{c[i],t[i]},\sigma_{i}\sim N(\mathrm{logit}(\phi_{c[i],t[i]}),\sigma_{i}^{2}),$ (7)
with $\sigma_{i}^{2}=s_{i}^{2}+\mathbb{I}(d[i]\in\mathcal{D})\sigma_{i}^{(\mathsf{source})2}+\mathbb{I}(i\in\mathcal{C})\sigma_{i}^{(\mathsf{char})2}+\mathbb{I}(i\in\mathcal{I})\sigma_{i}^{(\mathsf{outlier})2}$. We assume observations are conditionally independent except for data collected by the PMA program, as results from consecutive PMA surveys may be correlated because of repeated sampling of individuals across survey rounds. For PMA data series for a country-marital group, we use the multivariate normal specification:
$\bm{z}|\bm{\phi},\bm{\Sigma}\sim N(\mathrm{logit}(\bm{\phi}),\bm{\Sigma}),$ (8)
where $\bm{z}$ refers to the vector of logit-transformed observations, $\mathrm{logit}(\bm{\phi})$ refers to the respective mean vector of $\bm{z}$, and $\bm{\Sigma}$ to its associated covariance matrix. We set $\Sigma_{j,l}=Cov(z_{j},z_{l})=\sigma_{j}\cdot\sigma_{l}\cdot\rho_{\mathsf{PMA}}^{|t[j]-t[l]|}$, where $\rho_{\mathsf{PMA}}$ refers to the autocorrelation in PMA data errors.
Illustrative findings
Survey data and FPET fits are shows for Burundi, Ethiopia, and Niger in Figure 6. The fit for Burundi shows that based on the data outlier error set up, the estimates smooth over the second-last data point. In Ethiopia, FPET estimates are close to the most recent DHS, accounting for correlation in the PMA data series, that trends lower. FPET estimates suggest some uncertainty in recent years in Niger given conflicting data from PMA and a national survey. In this instance, if there is information on data quality, users may label a more recent survey as non-outlying to produce estimates based on that type of input (see discussion).
Figure 6: Data and estimates for mCPR among married women for Burundi, Ethiopia, and Niger. Data are shown in colored dots with 95% confidence intervals that reflect total error variance. FPET estimates are shown in black, with grey shaded areas representing 90% credible intervals.
4.4 Data model for service statistics data
Service statistics data are used in FPET to inform mCPR estimates. There are four types of FP service statistics data: the number of contraceptive commodities distributed to clients, the number of contraceptive commodities distributed to facilities, the number of times clients interacted with a provider for contraceptive services, and the number of current contraceptive users of any method. Track20 developed a tool to calculate an Estimated Modern Use (EMU) indicator from these different types of service statistics data [Track20, 2020, 2024]. EMU estimates have been used in FPET to inform mCPR estimates in the absence of recent survey data [Cahill et al., 2018]. Here we summarize the approach used in the most recent version of FPET. Additional details can be found in Mooney et al. [2024a].
To reduce risk of bias, we use observed EMU rates of change to inform mCPR estimates, as opposed to EMU level estimates. The motivation for using rates of change comes from difficulties capturing some contraceptive methods in specific EMUs or country settings. For example, in some countries, EMU estimates of IUD use based on service statistics data do not align with what is expected from survey data. In other settings, some methods cannot be adequately captured, such as sterilization when using commodities supplied to clients data, or condom use. When methods are missing, EMUs are biased but changes in EMUs, i.e., increases or decreases, are not necessarily biased. We note that percentage point changes in mCPR are the sum of method-specific percentage point changes, hence observed changes in mCPR do not depend on methods that do not change with time. This explains the rationale for using rates of change in EMUs in FPET: the risk of bias is lower for rates of change as compared to levels in settings where some methods may be difficult/impossible to capture in service statistics data.
When including rates of change of EMUs in FPET, we account for observation-specific uncertainty derived during EMU calculation, as well as uncertainty based on a country-type specific assessment of error. The data model is given by:
$\Delta EMU_{d,c,t}|\sigma_{EMU,d,c,t}\sim N(\Delta\mathrm{logit}(\phi_{c,t}),\sigma_{EMU,d,c,t}^{2}),$ (9)
where $\Delta EMU_{d,c,t}$ refers to the observed change in EMU for data type $d$, country $c$, and year $t$. Error variance $\sigma_{EMU,d,c,t}^{2}$ is given by
$\sigma_{EMU,d,c,t}^{2}=s_{EMU,d,c,t}^{2}+\sigma_{EMU,d,c}^{2}$ (10)
with $s_{EMU,d,c,t}^{2}$ the pre-calculated variance based on uncertainty in inputs and adjustments, and $\sigma_{EMU,d,c}^{2}$ a country-type specific variance. We obtain $s_{EMU,d,c,t}^{2}$ based on a Monte Carlo simulation, using uncertainty in inputs associated with the adjustments of EMUs to capture the private share (Mooney et al. [2024b]). The country-type variance is estimated hierarchically:
$\log(\sigma_{EMU,d,c})|\mu_{EMU,d},\sigma_{EMU}\sim N(\mu_{EMU,d},\sigma_{EMU}^{2}),$ (11)
with type-specific mean $\mu_{EMU,d}$, and overall variance $\sigma_{EMU}^{2}$, both of which are hyperparameters, estimated hierarchically using training data from all countries. Through this set-up, the country-type variance is based on global type-specific variance, as well as country data on the difference between change in EMU and survey-informed change in mCPR. As a result, the total variance associated with an EMU-based rate of change will be smaller for countries in which EMUs are more closely tracking survey-based estimates of mCPR.
Illustrative findings
The effect of including EMUs is illustrated in a case study in Figure 7. The top row shows information on mCPR (left) and rates of change in mCPR (right), based on EMUs (red) and survey-based estimates of mCPR (black). The vertical lines indicate the 95% confidence intervals for mCPR based on individual EMUs and their respective $s_{EMU,d,c,t}$ alone. The EMU-based rates of change suggest a stall after the most recent survey year, with rates of change that are lower than survey-based predictions. Plots in the bottom row illustrate the effect of including the EMUs in FPET estimates: the stall in EMUs past the most recent survey results in a slower estimated mCPR growth, as compared to survey-only estimates.
Figure 7: Illustration of the effect of including EMUs when estimating mCPR. The top row shows observed EMUs (red) for the level (left) and rate of change (right). The vertical lines indicate the 95% confidence intervals based on $s_{EMU,d,c,t}$ alone. Survey-based estimates of mCPR are added in black. The blue dashed line indicates the most recent survey year. The top row illustrates that the EMU-based rates of change suggest a stall after the most recent survey year, with rates of change that are lower than survey-based predictions. Plots in bottom row illustrate the effect of including the EMUs in FPET estimates: the stall in EMUs past the most recent survey results in a slower mCPR growth, as compared to survey-only estimates.
4.5 Computation and sequential model fitting
The model was implemented in the Stan programming language [Stan Development Team, 2024, Gabry et al., 2024], and all analyses were conducted using the R statistical computing environment [R Core Team, 2022]. Code is available in an open-source R package [Alkema et al., 2024b].
In a flexible process model that combines short- and long-term trends, it may be challenging to identify short-term acceleration or deceleration versus a longer term change in rate of change. This challenge holds true for the transition model used in FPET: rates of change combine longer term trends, captured by the transition function, with shorter-term accelerations or decelerations that are captured by the smoothing term. To overcome possible issues, we carried out sequential fitting to first estimate long-terms trends, followed by the assessment of short-term fluctuations. To estimate long-term trends, we fit a model with transition functions without smoothing components to the global data base. This step results in the estimation of subcluster transition functions that represent long-term trends. We use the estimated subcluster transition functions from this step as inputs in subsequent model fitting steps, i.e., to estimate parameters that govern short-term trends, as well as data model parameters.
The main utility of FPET is as a tool for countries to produce estimates for a particular population of interest. We provide this implementation through fixing non-country-specific parameters at point estimates from the global fit. The resulting “local model” can be fit to data for just one country. A web-based application of the local model with a user-friendly interface is available at https://fpet.track20.org/. This tool allows for user inputs in instances where substantive knowledge conflicts with automated categorization of observations into being possibly outlying or not. Specifically, if an observation is outlying but is deemed to capture the truth, the user has the option to change the possible outlier status and hence remove the outlier error term from the data model. Vice versa, when an observation from the reference category is deemed subject to data quality concerns, the error term for possible outliers can be added.
4.6 Model validation
Model development and checking took place in different stages, based on research on transition models and data models. We carried out a validation exercise in which we fit FPET to a training data set that excluded survey data that had been collected in or after 2018. The goal was to check how well the estimates and predictions based on the training set, i.e. estimates and forecasts that would have been produced in 2018, aligned with: (1) the estimates based on the full data set; and (2) the held-out survey data reported for years after 2018. We calculated summary measures based on prediction errors and calibration.
Detailed findings are in the appendix Section 8.2. In summary, we find that FPET is reasonably well calibrated for predicting mCPR. The point predictions that would have been constructed in 2018 generally fall within the uncertainty intervals constructed using all data at nominal levels or above. Larger absolute errors occur in settings where predictions are off due to an unforeseen stall or acceleration, or where recent data suggests an erroneous trend. In such instances, predictive performance may improve when including EMUs (Mooney et al. [2024a]).
5 Discussion
In this paper, we described the Family Planning Estimation Tool and the interplay between statistical modeling and usage in practice. We now turn to discussing takeaways and lessons learned from FPET that we believe are relevant to statistical modeling for demographic and global health indicators more generally. We end with some open challenges that also generalize to the broader field.
5.1 Takeaways and lessons learned
5.1.1 Collaborations between users and modelers are a key to success
Based on our experience with statistical modeling for family planning, as well as other reproductive, maternal, and child health indicators, we have concluded that the most important (but often overlooked) modeling challenge is to produce results that are reasonable for all populations and settings. An 80-20 rule in the context of statistical modeling for global health indicators would be as follows: producing reasonable estimates and forecasts for 80% of population-periods is easy; the remaining 20% is hard. While an easy-to-build-low-on-assumptions-type model based on, for example, commonly used generic spatio-temporal smoothing methods, may predict well when considering average results across populations and settings, it typically produces unrealistic results in a subset of settings. Moreover, the 20% of population-periods where models may fail to produce appropriate estimates are often those settings where the estimation matters most, as data are typically sparse or uncertain, or the outlying-ness of outcomes is something that needs to be reflected in the model-based estimates.
To obtain realistic model-based estimates for all settings of interest, including the 20% of population-periods that are unusual in some sense or where data are limited, models need to be built by a collaboration of statisticians and subject experts. Described in a less formal way (see Figure 8): to obtain realistic model-based estimates for all settings of interest, models need to be built by a collaboration between “data crunching” Bugs Bunnies and “this is how the world works” Brainy Smurfs. While the Bugs Bunnies can “easily” introduce powerful statistical tools, the Brainy Smurfs are able to provide reality checks regarding appropriateness of assumptions made and resulting findings. Indeed, the evolution of FPET, in terms of its set-up and assumptions made, was driven by a collaboration of data crunchers and substantive experts through the Track20 project. Active usage and feedback on modeled results resulted in the various updates that we described in earlier sections.
Figure 8: Bugs Bunny, Brainy Smurf, and Bayes. To obtain realistic model-based estimates for all settings of interest, models need to be built by a collaboration between “data crunching” Bugs Bunny and “this is how the world works” Brainy Smurf. The Bayesian framework can be considered for producing model-based findings.
5.1.2 Local tools are another key to success
We believe that local tools are incredibly important for successfully producing locally-relevant estimates. At a minimum, these tools allow for production of estimates based on data from a particular setting alone. It is typically easier to get buy-in for a set of estimates that is produced in-house, based on a fit to data that has been vetted, as opposed to estimates that have been produced in a top-down approach.
But the use and usefulness of local models go well beyond the production of just one set of estimates. Having a model that can be run annually builds a mechanism for data review, of surveys but most importantly service statistics. Track20’s experience suggests that countries have improved service statistics over the project because they are reviewing them annually, usually in a consensus meeting with many stakeholders, and it prompts follow-up action. Local tools also enable sensitivity analyses of various forms: models can be fitted to specific data sets or based on changing model assumptions. These kinds of fits can be helpful for analysts to better understand the data and model. Track20 has successfully used this approach in consensus meetings, as well as in additional technical meetings with various government and non-government stakeholders. For example, in India and Kenya, the tool has been used to analyze data outliers and inform decisions related to data exclusion.
In summary, by enabling in-house production of estimates and additional assessments, local tools empower governments and move away from a commonly used approach of externally produced global health estimates.
5.1.3 Communication and training needs to be geared at different audiences
Given the range of audiences involved in modeling exercises, clearly communication needs to be geared at these different audiences and, depending on the audience, vary in terms of technical detail. To make this possible, communication of model assumptions and explanations of “how the model works” are best produced by different actors as well. In our case, the statisticians produced technical writings while Track20 focused on producing materials and supportive documentation that was targeted to less statistical audiences. The local tool can also be helpful for communication, e.g. through sensitivity analyses that illustrate how the model responds when fitted to different sets of data.
5.1.4 Go Bayesian!
We have found that the use of Bayesian models enabled the development and evolution of FPET. The Bayesian framework provides flexibility in translating assumptions into probabilistic models. The wide availability of Bayesian modeling software facilitates the easy implementation of such models and active development of computational approaches have resulted in reduced computation times and expanded the types of models that can be fitted. Last but not least, the Bayesian paradigm also allows for uncertainty to be accounted for in, and propagated across various model components. In our set-up, we derived a local tool from a global model, as a pragmatic, solution-focused approach. The Bayesian framework facilitates this type of set up.
5.2 What’s next: Open challenges
5.2.1 How to allow for context-specific adjustments to model-based estimates?
The use of model-based estimates can be a double-edged sword: On the one hand, model-based estimates can provide insights but on the other hand, model-based estimates may be based on assumptions that do not hold true or estimates may be subject to substantial levels of uncertainty, limiting their usefulness for action. Moreover, as mentioned in relation to the 80-20 rule, the population-periods where models may fail to produce appropriate estimates are often those settings where the estimation matters most, as data are typically sparse or uncertain, or the outlying-ness of outcomes is something that needs to be reflected in the model-based estimates. It is worth noting that the increased focus on capturing within-country differentials compounds these issues: the interest in smaller subgroups brings about greater data needs and – when data availability does not match these needs – exacerbates the importance as well as the potential limitations of models.
When dealing with a setting where model-based estimates conflict with substantive knowledge, i.e. when model-based estimates do not reflect the truth, what can a user or modeler do? Examples from FPET fits are settings where either estimates smooth over a data point (the model considers it an outlier) while expert knowledge indicates that the data point captures the situation very well, or vice versa, where data of dubious quality or from populations that may differ from the population of interest have a great influence on the estimates. In earlier work, we came up with some approaches to deal with some of these settings. For example, users can indicate whether data points refer to populations that may not be representative of the population of interest, and the model fit takes that into account [Alkema et al., 2013]. In the currently-used FPET data models, users can make updates to the default categorization of observations into the “possible outlying” category, to deal with instances where substantive knowledge conflicts with automated categorization of observations into being outlying or not. These FPET approaches are examples of set-ups where users can provide input to improve upon estimates.
While giving the user options to provide additional context related to data or the outcome of interest is a positive way forward to producing high-quality estimates, the open challenge relates to what information to ask for, and how that information can be used. While some context-specific questions may have reasonably clear answers that can be used in modeling, others may not. For example, when considering data quality and errors, it is likely easier to answer the question whether a survey is representative of the entire country or limited to specific regions as opposed to a question related to data quality more generally. In addition to difficulties obtaining relevant information, another challenge is related to the need for standardization of estimation procedures and keeping this process as objective as possible. The option of user inputs should not enable manual tweaking of estimates to optimize gains that are not statistical accuracy but instead, could relate to political agendas. Given these challenges, the process of obtaining context-specific information (and documentation) and its usage needs to be carefully considered.
5.2.2 How to improve the process of model development and communication?
Model development and communication of strengths and weaknesses of model-based estimates is an area of great importance. We believe that the community of statistical modelers have a key role to play in improving on the status quo. The current literature on statistical modeling of global health indicators lacks standardization of the specification and communication of model assumptions. Because of this gap, we find that models are difficult to compare and modeling assumptions may be difficult to understand. In addition, guidance to systematize model building is very limited. This may result in model choice being based on prior experiences of the research team involved as opposed to suitability considerations.
To advance the development and use of model-based estimates, we see a need to develop material to help standardize the specification and communication of model assumptions. To work towards that goal, we proposed the use of a new model class, referred to as Temporal Models for Multiple Populations (TMMP, see Susmann et al. [2022]). This model class facilitates documentation of model assumptions in a standardized way and facilitates comparison across models. We show how existing models for a variety of indicators can be written as TMMPs and how the TMMP-based description can be used to compare and contrast model assumptions. This approach, or other efforts, can be considered to help communicate model assumptions.
There are opportunities to move toward systematizing model building. Given the typical challenges associated with the 80-20 rule, i.e., getting the 20% right, it does not seem realistic to aim for a standard recipe for model building that can be applied across a range of indicators and settings and research questions. However, we do see an opportunity to consider the process of model building in a more systematic manner, through ideas that have been recently put forward under the Bayesian workflow umbrella. In our recent work on FPET, we have introduced elements of this type of workflow that are relevant to our modeling goals. Last but not least, modelers cannot (and should not aim to) pretend that model-based estimates have the final word in settings where additional data are needed to answer open questions. A more targeted approach to data collection, one that is informed by statistical models, could result in great gains in insights.
5.2.3 How to increase diversity in the modeling community?
To maximize the value of statistical models for demographic monitoring, action is needed to expand and diversify the community of actors involved, ranging from end users of model outputs to advanced users (applied statisticians and data scientists who understand the detailed workings of models), and modelers who develop and update models.
To date, the great majority of major modeling efforts are based on a set up where academics or international organizations based in the global north develop models and often also produce estimates. The Track20 project with FPET can serve as an example of how to diversify this community, through local capacity building that enabled local usage of tools and local production and ownership of estimates, and through improved interaction between local users and other actors involved. Additional opportunities are to diversify communities of modelers and advanced users so that there are more statistical models being produced in the global south. The expansion of the community of advanced users is especially useful when thinking about rolling out tools that have more advanced usage options.
The opportunities to diversify communities of modelers and advanced users are contingent upon outreach to, and capacity building of under-represented groups, in particularly based on geography. The good news is that with the availability of high-quality open-source software tools such as R and Stan, and the increased availability and quality of online training material in statistics and data science in various forms, access and training are no longer limited to audiences that are geographically close.
With the combination of initiatives like Track20, that diversify the community involved in model-based estimates on specific areas, and the opportunities provided through open-source software and expanded communities of individuals trained in statistics and data science, a more diverse modeling community could become reality in the short-term.
6 Conclusions
Using our experience with the Family Planning Estimation Tool, we discussed take aways and lessons learned related to the field of statistical modeling for monitoring for demographic and global health indicators. We discussed that (1) Interaction between modelers and users/local experts is essential to producing high quality model-based estimates; (2) Local tools are a key to success; (3) Communication and training needs to be geared at different audiences; and (4) Going Bayesian allows for modeling flexibility. We ended with a more general call to action to the community of statisticians and statistical demographers in the form of steps to consider, to optimize the relevance of statistical modeling in practice. Our call to action focuses on open challenges: (1) How to allow for context-specific model adjustments, for improving estimates in specific settings, (2) How to improve the process of model development and communication, and (3) How to diversify the modeling community. While challenges remain, so do continued advancements in the field of statistics and data science, improvements in access and training opportunities, and the need for high quality measurements. Given all that, we are very optimistic about the future of field of statistical modeling for monitoring for demographic and global health indicators.
7 Data availability
Data and code are available at https://github.com/AlkemaLab/fpet.